Kermit is considering purchasing a new computer system. The purchase price is $128663. Kermit will borrow one-fourth of the purchase price from a bank at 10 percent per year compounded annually. The loan is to be repaid using equal annual payments over a 3-year period. The computer system is expected to last 5 years and has a salvage value of $6897 at that time. Over the 5-year period, Kermit expects to pay a technician $20,000 per year to maintain the system but will save $76807 per year through increased efficiencies. Kermit uses a MARR of 12 percent to evaluate investments. What is the net present worth for this new computer system? 


Question 8

Consider palletizer at a bottling plant that has a first cost of $135893, has operating and maintenance costs of $15800 per year, and an estimated net salvage value of $28768 at the end of 30 years. Assume an interest rate of 8%, what is the future worth of this project?
Answer #1
Purchase Price              1,28,663
Loan (1/4 of Purchase Price)                 32,166
Int 10% Compounded annually
Loan time period 3
Salvage Value             6,897.00
Useful Life 5
Annual Maintenance Cost           20,000.00
Savings due to increased efficiencies 76807
MARR 12%
Loan Calculation
X Principal           32,165.75 Formula = (128663-32166)
Y Interest at 10% for 3 years           10,646.86 Formula = (1*(1+10/100)^3)
Total Loan to be Paid (X+Y)           42,812.61
Outflows
Time period (t) 1 2 3 4 5
A Initial Investment (Purchase Price - Loan Value)             96,497.3
B Loan EMI             14,270.9    14,270.9    14,270.9
C Annual Maintenance Cost 20000 20000 20000 20000 20000
D Total Outflows (A+B)              1,30,768        34,271        34,271        20,000        20,000
E Present value Factor @ 12% (1/(1+12/100)^t)                     0.89            0.80            0.71            0.64            0.57
F Present Value of Outflows (D*E)              1,16,757        27,321        24,393        12,710        11,349
Inflows
G Annual Savings due to increased efficiencies                 76,807        76,807        76,807        76,807        76,807
H Salvage Value          6,897
I Total Inflows (G+H)                 76,807        76,807        76,807        76,807        83,704
J Present value Factor @ 12% (1/(1+12/100)^t)                     0.89            0.80            0.71            0.64            0.57
K Present value of Inflows (I*J)                 68,578        61,230        54,670        48,812        47,496
L Total Present value of Outflows              1,92,530
M Total Present value of Inflows              2,80,786
NPV (M-L)                 88,256

Question 8

Initial Cost        1,35,893.00
Annual Maintenance Cost           15,800.00
Salvage Value at end of 30 Years           28,768.00
Interest rate % (R ) 8
Future Worth
Time period (t) 30
Initial Cost              1,35,893
A Future value after 30 Years           13,67,445 Formula = 135893*(1+(1+R/100)^t)
B Future value of Annual Maintenance Cost              1,58,990 Formula = 15800*(1+(1+R/100)^t)
C Salvage Value at end of 30 Years 28768
Value of the Project after 30 years (A-B+C)           12,37,223